What are the dimensions of a rectangular box that has w=x h=x and length=x+1? I have given up! I came up with answers that did not match, and then tried to put answers into the quadratic formula, and still nothing. Your help would be much appreciated. Thank you.
Answer by stanbon(75887)

What are the dimensions of a rectangular box that
has w=x h=x and length=x+1? total Surface Area of the box = 130cm^2.
------------------
Equation:
SA = 2(lw)+2(lh)+2(wh) = 2(lw+lh+wh)
---
130 = 2[(x+1)x + (x+1)x + x^2]
65 = x^2+x+x^2+x + x^2
3x^2 + 2x-65 = 0
----
Use the quadratic formula to get:
x = 4.333
-----
width = 4.333
height = 4.333
length = 5.333
=====================
Cheers,
Stan H.
=================